### v3

# Failure Modes of LLMs for Causal Reasoning on Narratives

Abstract
The ability to robustly identify causal relationships is essential for autonomous decision-making and adaptation to novel scenarios. However, accurately inferring causal structure requires integrating both world knowledge and abstract logical reasoning. In this work, we investigate the interaction between these two capabilities through the representative task of causal reasoning over narratives. Through controlled synthetic, semi-synthetic and real-world experiments, we find that state-of-the-art large language models (LLMs) often rely on superficial heuristics—for example, inferring causality from event order or recalling memorized world knowledge without attending to context. Furthermore, we show that simple reformulations of the task can elicit more robust reasoning behavior. Our evaluation spans a range of causal structures, from linear chains to complex graphs involving colliders and forks. These findings uncover systematic patterns in how LLMs perform causal reasoning and lay the groundwork for developing methods that better align LLM behavior with principled causal inference.
1 Introduction
Large Language Models (LLMs) have achieved promising performance across a range of tasks, as a result of their ability to absorb rich world knowledge from large-scale unsupervised data. As such, there is rising interest in deploying them in autonomous and agentic settings. However, a key prerequisite for autonomous decision making is causal reasoning: agents must be able to move beyond relying on associations and be capable of robustly determining the consequences of their actions. Does this capability arise naturally from large-scale pretraining?
Robust causal reasoning is particularly challenging as it relies on a combination of knowledge and reasoning capabilities. Unlike mathematical reasoning benchmarks (Cobbe et al., 2021) – which draw from a relatively constrained set of problem solving strategies and results – arriving at correct causal inferences often requires leveraging domain-specific knowledge about the events or variables involved. On the other hand, models must go beyond blindly retrieving memorized associations and knowledge to identify the correct relationships under atypical or counter-intuitive settings. World knowledge and reasoning must be carefully balanced for robust causal reasoning: incorporating some memorized world knowledge can be helpful, but over-reliance can lead to errors in atypical settings.
Prior works have primarily studied reasoning and world-knowledge of LLMs separately. For example, benchmarks on mathematical reasoning or coding typically study these capabilities in isolation – with minimal external world knowledge needed to solve problems. On the other hand, benchmarks for knowledge intensive tasks can generally be solved by simple retrieval of memorized knowledge. Thus, the interplay of knowledge retrieval and reasoning (and potential conflicts between them) remains understudied.
In this work, we study this interplay in the setting of causal reasoning over textual narratives. We pay particular attention to (a) long sequences of events (as are likely to arise in autonomous LLM settings) and (b) atypical settings where typical “common-sense" causal relationships may not hold. Concretely, we start by considering settings where there is an (unknown) underlying causal chain graph of the form $V_{1}\rightarrow V_{2}\rightarrow\ldots\rightarrow V_{N}$, where each node $V_{i}$ has some semantic meaning (e.g., smoking), that is verbalized in the form of a (realistic) narrative $N$. For a given narrative $N$, we consider two causal reasoning tasks: (1) Does $V_{i}$ have a causal effect on $V_{j}$? and (2) Given the node identities $(V_{1},\ldots V_{N})$, construct a causal chain graph that is faithful to the narrative. While these tasks do not encompass all aspects of causal reasoning (e.g., counterfactual reasoning), they are important primitives for successful causal reasoning. We then extend our analysis to more complex underlying graph structures including colliders and forks.
Our primary contribution is to characterize when large language models are able to robustly reason under these causal structures. On the one hand, we find that LLMs display distinctive failure modes related to interference between reasoning and world knowledge in causal inference. Firstly, we show that LLMs are influenced heavily by a prior that causes are likely to appear before effects in a narrative. We observe that when the narrative is constructed in the reverse topological order of the causal chain (i.e., the edge $V_{i}\rightarrow V_{i+1}$ is narrated before $V_{i-1}\rightarrow V_{i}$), the performance of the LLM suffers as it often assigns the cause to an earlier event and the effect to a later event in the narrative. Secondly, we show that LLMs use their parametric causal knowledge (i.e., if an event typically causes another event) as a shortcut to answer causal questions. Thus, when the cause-and-effect pairs implied by the narrative conflict with the parametric knowledge, the LLM often ignores the specifics of the narrative and defaults to its parametric knowledge. Neither prompting with Chain of Thought (CoT) (Wei et al., 2022) nor In-Context Learning alleviates these failures.
However, LLMs are much less impacted by variation in the reasoning difficulty of the task when the prompting scheme explicitly isolates reasoning and world knowledge. First, we find that asking the LLM to extract the entire causal graph implied by the narrative results in a high degree of success at correctly ordering individual events, largely avoiding both failure modes described above. However, these benefits dissipate if the model is prompted to use the extracted graph alongside the narrative. Second, LLMs exhibit only slight performance degradation when reasoning over narratives that display more complex graph structures than chains, for example forks or colliders. Third, while LLMs often struggle with longer narratives containing more events, this failure is also substantially mitigated by asking the LLM to just extract a graph. All together, our results paint a more nuanced picture of LLMs’ causal capabilities than simple success or failure and suggest that future development should focus on isolating and then composing LLMs’ strengths at reasoning and world knowledge in order to avoid conflicts between them. We validate our findings across both controlled synthetic (LLM generated) narratives and real-world articles from CauseNet Heindorf et al. (2020).
Ultimately, our work presents a comprehensive study of causal reasoning in LLMs. Our findings reveal that current state-of-the-art models can fail to reliably perform simple causal reasoning tasks on narratives. We distill two shortcuts that underlie this unreliability: over reliance on positional biases and parametric knowledge. On the other hand, we find that incorporating an explicit causal graph identification task can elicit significantly better reasoning behavior.
2 Related Works
Causal Reasoning in Large Language Models
Jin et al. (2023) develop a benchmark for testing causal reasoning in LLMs given causal graphs, finding that language models can struggle with the task. However, the queries examined in Jin et al. (2023) require probability calculations, potentially conflating causal reasoning and arithmetic failures. Tan et al. (2022) shows the capability of a neural network trained on news data to label causal structures in individual sentences. Joshi et al. (2024a) chronicles failure modes in textual, but non-narrative form data (e.g. text formulaically written as Event 1 Causes Event 2 Causes Event 3 Causes Event 4). Our paper expands upon such a line of work by testing the LLM’s abilities in both real and synthetic texts that would plausibly be seen in everyday life. Another contrasting work, Jin et al. (2024), uses only statistical language indicating event correlations as input.
(Gordon et al., 2012; Joshi et al., 2024b; Ho et al., 2023; Zhang et al., 2023; Wang et al., 2023; Ashwani et al., 2024) study causal reasoning ability as it relates to inferring causal relations based on "common sense". In such common-sense based settings, it is straightforward for models to simply rely on memorized knowledge from pretraining and achieve good performance, without leveraging any more general causal reasoning capabilities. Our work seeks to disentangle this general causal reasoning ability by specifically testing cases where causal relationships may contradict common-sense knowledge. This serves as a more robust measurement of the causal reasoning capabilities in unfamiliar and atypical scenarios. Empirically, we show that models struggle significantly in adapting to unfamiliar causal relations.
Another important distinction of our work is the focus on longer-form narratives. Existing works such as (Gordon et al., 2012; Zečević et al., 2023; Ho et al., 2023; Frohberg and Binder, 2022; Li et al., 2023; Gao et al., 2023) primarily examine short-form questions about a single causal relationship. On the other hand, our work examines longer and more complex sequences of events. Moreover, in contrast to domain-specific question banks such as Intuitive Physics studied in (Zečević et al., 2023), our narratives examine a more diverse range of topics (as illustrated by the sample narratives presented). As a result, our dataset provides a more realistic and diverse examination of LLM causal reasoning capabilities than prior works. As such, our work is unique in that we are the first paper to analyze non-common sense based causal reasoning in narratives that use everyday language.
Causal Story Generation
Kıcıman et al. (2024) shows that LLMs have strong abilities to generate causal texts. Ammanabrolu et al. (2020) introduces soft causal relations—causal constraints that match what readers expect—and uses commonsense inferences to bridge high-level plot points, resulting in more coherent narratives that align with everyday causal expectations. Tian et al. (2021) contributes by employing counterfactual knowledge to generate hyperboles, making story generation more realistic. Li et al. (2022) shows that asking a model to explain a cause or effect by generating new text conflates language generation with prediction; instead, their approach asks the model to simply indicate the sentence number representing the cause or effect, leading to stories that better respect causal relations. For our synthetic text generation, we focus on creating narratives that are extremely explicit and simple. In contrast to Ammanabrolu et al. (2020) and Li et al. (2022)’s approach of bridging events using commonsense, our narrative scheme already embeds explicit causal language between events that are causally related so that no inference or common-sense reasoning is required from the reader to reason about causality. Furthermore, our experiments often intentionally contradict common-sense parametric knowledge to check the model’s ability to solely rely on the self-contained narrative. Similarly, regarding Tian et al. (2021), we opted to avoid abstract language structures that might confuse even human readers.
3 Experiments with Synthetic Data
3.1 Setting
Synthetic Narrative Generation
In our synthetic experiments, we use three leading LLMs: OpenAI’s GPT-4o (OpenAI et al., 2024), Anthropic’s Claude 3.5 Sonnet (Anthropic, 2024), and the open source LLama 3.1 8b (Grattafiori et al., 2024). While we focus on GPT-4o in the main text, results from other models are in the Appendix. The purpose of our synthetic setup is to carefully control the conditions under which the LLMs are tested. In terms of the general setup of our fully synthetic experiments, we first use the LLM to generate events (which are real world phenomena like rain or plants growing). Then these events are linked together into a chain graph $G$ that acts as the causal ground truth (eg rain $\rightarrow$ plants growing). The LLM is given $G$ and asked to create a narrative that stays faithful to the causal relationships in $G$. These narratives are checked by researchers to ensure consistency with their base causal graphs. More specifically, when constructing the dataset, we asked researchers (select authors who were blind to the true underlying graph) to reconstruct the causal chains given just the narratives, and 98 percent of the time (out of 100 random samples), the humans were able to find the unique correct causal ordering. Roughly 2500 narrative samples were generated. To ensure a variety of events go into the narratives, we generate 100 to 1000 distinct events at a time and randomly pick the small number needed for narrative construction (all narratives are in supplementary files and select narratives are in the Appendix).
Providing only the narrative as input (and not $G$), we then ask the LLM to find $G^{\prime}$, the predicted underlying causal structure expressed by the narrative. In other words, the LLM is asked to output a causal graph that it thinks embodies the relationships in the narrative. Next, a series of causal questions is created by randomly sampling $10$ tuples of events from $G$ and asking the LLM whether an event in the tuple causes the other based on the narrative and/or $G^{\prime}$.
Prompting Strategies
We evaluate five prompting styles for causal reasoning where the names in italics represent those used in the legends of figures: Standard QA Prompting (Standard), where the model is simply asked to identify the causal relation between two narrative events; Chain-of-Thought (CoT), which instructs the model to articulate step-by-step reasoning before answering; In-Context Learning (In-Context), which precedes the query with illustrative question–answer examples; Explicit Causal Graph Extraction (Graph), which asks the model to generate an entire causal graph $G^{\prime}$ over all events and assesses whether the ordering of the target pair is correct; Narrative-Augmented Graph Extraction (Narr-Graph), which first elicits $G^{\prime}$ and then supplies both $G^{\prime}$ and the original narrative for joint reasoning about the causal pair. Exact prompts are in Appendix A.
[…]
3.2 Impact of Event Ordering
Our experiments show that LLMs rely on the ordering in which the events are verbalized in a narrative when determining causal relationships. To investigate this, we started with randomly generated events that were used to make a ground truth graph $G$. During the creation of the narrative, we specified that the LLM either places the events in (1) the order that matches the topological causal ordering of the graph (e.g., if event $A$ (indirectly or directly) causes $B$, then event $A$ is mentioned before $B$ in the narrative), or (2) a way that runs opposite to the causal ordering (event $B$ would be mentioned before event $A$ in the narrative even though $A$ (directly or indirectly) causes $B$). We refer to these as the Forward and Reverse topological ordering, respectively. As an example, the following is a GPT-4o generated Reverse topological narrative for the causal chain: Art $\rightarrow$ Wine tasting $\rightarrow$ Charity fundraiser:
The charity fundraiser was made possible because of the successful wine tasting event that attracted numerous generous patrons. The wine tasting was organized as a result of the art exhibition drawing in a sophisticated audience interested in cultural experiences.
Each edge in the narrative is verbalized in the opposite order to its place in the causal chain. All narratives can be found in the linked code.
LLMs Rely on Event Ordering Across Prompting Strategies
As shown in Figure 1 (left), in the Forward direction, standard QA, CoT, and In-Context prompts all perform very well. This is in contrast to the Reverse orientation when we look at the performance of the standard QA, COT, and In-Context prompts. From this plot, we can see that naive COT and In-Context prompting do not seem to significantly boost accuracy under our conditions. Perhaps more interestingly, we find that the way the LLM answers questions using the narratives is not always consistent with the causal graph $G^{\prime}$ that the LLM builds when asked to predict the underlying graph structure (see consistency plot in right side of Figure 1, where consistency measures agreement between the answers of the LLM and $G^{\prime}$). In the Reverse orientation, answers given by the extracted causal graph G’ and the previously discussed prompting strategies seem to differ greatly. Additionally, the trend of those prompting strategies on the consistency plot for the Forward orientation narratives (comparing performance to $G^{\prime}$) mirrors their trend on the accuracy plot which compares performance to ground truth $G$ (left side).
Explicit Causal Graph Extraction Avoids Shortcuts
This led us to test the accuracy of only using the extracted graph $G^{\prime}$ to answer causal questions. In this case, once $G^{\prime}$ is extracted by the LLM, it is not given to the LLM again to answer questions (but rather used directly). We found that this strategy did significantly better in the Reverse direction than the other prompting strategies. Surprisingly, using $G^{\prime}$ in the Reverse direction narratives to answer causal questions did as well as using $G^{\prime}$ in the Forward direction narratives. Next, we tried prompting using the narrative and $G^{\prime}$ (the LLM is given $G^{\prime}$ in this case in the prompt). This technique could be thought of as a type of CoT prompting strategy. However, in the Reverse direction narratives, the increase in accuracy achieved by only using $G^{\prime}$ completely dissipates. We conjecture that the process of building the extracted Causal Graph $G^{\prime}$ forces the LLM to engage in long term reasoning instead of using the simple shortcut, but when the narrative is again provided - the LLM defaults back to the shortcut.
[…]
3.3 Impact of Parametric Knowledge (In)consistency
Experimental Setup
We also find that LLMs tend to rely on parametric knowledge when it is present, and can fail when narratives are inconsistent with the LLM’s parametric knowledge. To test this, we elicit the LLM’s pre-existing parametric knowledge when generating the event chains. We prompt the LLM to pick a series of events such that each event has some relation to the subsequent event – either the event is Causal to the next event (e.g., disease causes shorter lives) or the event is Anti-Causal (e.g., disease causes longer lives). For example, we might know that node $1$ is Anti-Causal to node $2$ from parametric knowledge. Thus, when we make the causal ground truth graph $1\rightarrow 3\rightarrow$ 2 (this disagrees with parametric knowledge), create a narrative from it, and then ask the LLM if node $1$ causes $2$ based on the narrative: it should say yes based on the narrative even though that disagrees with its parametric knowledge. After the ground truth graph is created, we generate the narrative in the Forward topological orientation to avoid confounding failure modes. The full process (along with illustration) explaining how the parametric and causal graphs are created is in Appendix A.2. As a textual example, assume that we know a parametric anti-causal link exists from stressful job to increased happiness, and from lack of sleep to improved cognitive function. We can then construct the causal chain Stressful JobLack of SleepIncreased HappinessImproved Cognitive Function. From this causal chain, we create the narrative:
The constant demands of a stressful job led to her experiencing chronic lack of sleep. Surprisingly, she found that the lack of sleep heightened her sense of euphoria, making her unusually cheerful at work. Increased happiness from this unexpected cheerfulness seemed to improve her cognitive function.
If the LLM is asked if a stressful job leads to increased happiness, the parametric knowledge shortcut indicates the answer should be no – however, the shortcut fails as the narrative indicates that a (indirect) causal link does exist.
Models Exploit Parametric Knowledge
We find that, in synthetic experiments, the LLM finds the correct causal relation generally only when that relation agrees with its parametric knowledge. This is exemplified in the plot in Figure 2 (left) where we see good performance on narratives that agree with parametric knowledge (Causal parametric knowledge) and poor performance on narratives that disagree with parametric knowledge (Anti-Causal parametric knowledge). We also notice an interesting phenomenon for the Anti-Causal parametric case where using just the extracted graph provides massive improvements over any prompting strategy that involves using the narrative to directly answer questions. This strategy is comparable in performance to when the parametric knowledge is Causal. It seems that the narrative may only serve to distract the LLM when parametric knowledge disagrees with the narrative.
3.4 Impact of Narrative Complexity
In the previous sections, we identified two shortcuts which models exploit in causal reasoning tasks. Here, we test the influence of narrative complexity on these failure modes. We examine two measures of complexity: (a) the narrative length and (b) the presence of complex graph structures.
Narrative Length
In conditions where the LLM exhibits failure modes (Reverse and Anti-Causal orientations), the performance also tends to decay as the size of the narrative and the number of events in the narrative increases. As we can see in Figures 1 and 2 (Left), it seems that the longer the narrative is, the more the LLM relies on shortcuts instead of performing reasoning. However, the extracted graph $G^{\prime}$ can often maintain a consistently high level of accuracy across narrative sizes even for cases when a failure mode would normally be exhibited.
Causal Graph Complexity
As the bulk of our work has focused on detecting the simplest failure modes possible, we studied narratives with an underlying chain graph structure. However, the presence of more complex causal structures in the narrative could exacerbate the existing failure modes or trigger novel failures. To study this, we create causal graphs utilizing two common causal structures: Forks (one node has a causal relationship to multiple other nodes) and Colliders (multiple nodes have a causal relationship to the same node). We generate narratives (the complete algorithm is described in Appendix A.3) such that each underlying causal graph contains at least one of these structures, and may randomly contain multiple such structures based on the size of the narrative. An example is shown in Figure 3.
Heavy RainfallPower OutageFlooded StreetsTraffic Jam
Narrative: The heavy rainfall not only caused a power outage in several neighborhoods but also led to flooded streets. The aftermath of the power outage (disabling traffic lights) and the flooded roads (blocking street access) caused a traffic jam.
Figure 3: Causal graph with story showing a fork (first sentence) and a collider (second sentence).
As can be seen in Figure 2 (right side), we find that while the LLM generally performs worse at reasoning about the complex narratives than simple narratives (with underlying chain graphs), the gap is very starkly less than can be seen in the other failure modes. This finding can be supported by (Dettki et al., 2025) which finds that GPT-4o reasons similarly to humans on a single sentence that describes one collider relation. Our work extends their work by using a long-form narrative based on a causal graph with potentially multiple colliders and forks instead of only one collider.
4 Experiments with Real World Causal Graphs
In this section, we extend our analysis to narratives involving real-world causal graphs from CauseNet (Heindorf et al., 2020), a large-scale knowledge graph of (claimed) causal relationships between real-world concepts. We perform experiments using the GPT-4o (OpenAI et al., 2024) and Llama-3.1 8B models for our experiments. We concentrate our analysis on the same factors (positional biases and parametric knowledge consistency) as explored in the semi-synthetic settings.
4.1 Experimental Setting
The CauseNet dataset can be represented as a collection of $D$ tuples $\{(C_{i},E_{i},\mathbf{S}_{i}\}_{i=1}^{D}$, where $C_{i}$ denotes the cause (e.g., fatigue), $E_{i}$ denotes the effect (e.g., accidents), and $\mathbf{S}_{i}$ is a set of sentences (extracted from Wikipedia and ClueWeb12 (Callan, 2012)) that entail a causal relationship from $C_{i}$ to $E_{i}$. We retrieve causal chain graphs $V_{1}\rightarrow V_{2}\rightarrow\ldots\rightarrow V_{N}$ of various lengths, where each causal relation $V_{i}\rightarrow V_{i+1}$ is from CauseNet and verbalize these chains as narratives in the following ways:
Semi-synthetic narratives.
In this setting, we use real causal graphs from CauseNet but synthetically verbalize them via the LLM. In particular, we prompt the LLM to generate sentences for each edge ($V_{i}\rightarrow V_{i+1}$) in the causal graph, while ensuring the sensibility of the entire narrative. For eammple, the following is a narrative for the chain fatigue $\rightarrow$ accidents $\rightarrow$ injury:
Fatigue can cloud judgment and slow reaction times, leading to an increase in accidents on the road. As a result, these accidents often lead to serious injury for those involved, highlighting the dangerous consequences of driving while fatigued.
Real-world narratives.
For the real-world narratives, the sentence for each edge is chosen from the CauseNet dataset. To ensure that the narrative as a whole remains coherent, we prompt the LLM to ensure that the sentences for every pair of adjacent edges logically follow each other. For example, the following is the narrative for the causal chain fatigue $\rightarrow$ accidents $\rightarrow$ injury:
Workers work long hours in mines and factories where fatigue and a lack of concentration can easily cause accidents. These accidents are the leading cause of injury in this country for people ages 1-34.
Additional examples of semi-synthetic and real-world narratives are presented in Appendix C.1 (the entire set of narratives used for our experiments is available in the linked code).
Prompting Strategies
For simplicity, we limit the prompting techniques used to (see Appendix C.2 for the prompt templates): Standard QA Prompting, Chain-of-Thought and Explicit Causal Graph Extraction. We evaluate the accuracy for each pair of nodes $(V_{i},V_{j})$ for the three prompting strategies on the semi-synthetic and real-world narratives.
4.2 Impact of Event Ordering and Chain Length
As described in the previous section, we verbalize each causal chain graph $V_{1}\rightarrow V_{2}\rightarrow\ldots\rightarrow V_{N}$ from CauseNet into a narrative in the forward and reverse topological order.
Figure 4: The accuracy of various prompting strategies (error bars denote 95% CIs). We observe that the accuracy is lower in the reverse direction (and tends to decay as the chains get longer).
In both the semi-synthetic (Fig. 4 left) and real-world narratives (Fig. 4 right), the Forward Graph strategy performs the best, with its accuracy remaining stable even as the chain length increases. We observe that Forward Standard and CoT outperforms Reverse Standard and CoT, with the Reverse accuracy declining substantially as the chain size gets large. We also see that in this regime, extracting the causal graph makes inference in the Reverse orientation competitive with inference in Forward.
4.3 Effect of Parametric Knowledge Consistency
Experiment Setup
Next, we analyze the extent to which the LLM relies on its parametric knowledge to answer causal reasoning queries as opposed to the causal structure expressed in the narrative. For every pair of nodes $(V_{i},V_{j})$ in the chain graphs, we elicit the parametric knowledge of the LLM by asking the LLM whether a causal effect between the two nodes would be atypical (see Appendix B.2 for the exact prompts utilized). Through these prompts, we identify cause and effect chains which contradict the model’s parametric knowledge. For example, in a chain graph from our dataset, there is a path from streambank erosion to higher prices, but this contradicts the LLM’s parametric knowledge since this causal effect may not typically exist in the real-world. In total, we find that roughly 5 percent of the relations in CauseNet violate the LLM’s pretraining knowledge. We sampled narratives from CauseNet until we got 100 (of chain sizes between 3 and 9) narratives that contain relations that violate the LLM’s pre-training knowledge and 100 that are consistent. These narratives are constructed in the Forward topological ordering to avoid confounding failure modes.
LLM Performance Suffers on Atypical Causal Relations
We evaluate the three prompting strategies separately on the subsets of cause-and-effect pairs that are in agreement and in conflict with the parametric knowledge (see Table 1). We observe that when there is no conflict (i.e., the parametric knowledge agrees with the causality expressed in the narrative), the accuracies with and without CoT are greater than $90\%$. However, when the parametric knowledge conflicts with the narrative’s causality, the accuracy is significantly lower, even with CoT. This suggests that when asked to reason about cause and effect in a narrative, the LLM seems to rely heavily on its parametric knowledge and is unable to grasp the specific causal chains expressed in the narrative itself (despite the causal chains as a whole being realistic).
Explicit Causal Graph Extraction Avoids Shortcuts
Interestingly, when using extracted graph for performing causal reasoning, the performance is very high, both with and without conflicts. This is likely because when asked to extract the graph from the narrative, the LLM pays more attention to the entire narrative as opposed to when directly queried on a cause-and-effect pair (where the LLM defaults to its parametric knowledge). These results show that even when the LLM constructs a reasonably good causal chain graph, the LLM does not leverage this graph when queried directly about the causal effects in the narrative (even with CoT), further highlighting the advantage of extracting the causal graph directly.
Standard CoT Graph
Semi-synthetic
Without Conflict 99.8 99.6 99.9
With Conflict 67.2 73.1 98.7
Real-world
Without Conflict 90.9 89.2 97.9
With Conflict 52.1 57.6 93.2
Table 1: The average accuracy across different narratives with the three prompting strategies partitioned by whether the cause-effect pairs conflict with the LLM’s parametric knowledge (we omit the 95% CIs as they are smaller than $0.3$).
[…]
4.4 Narrative Complexity
We can see from Figure 4 that LLM performance degrades with narrative length, especially when a failure mode is present. We furthermore experimented with complex narratives with causal graphs containing forks and colliders (full graph and narrative creation algorithm in Appendix B.3). We can see in Figure 5, that in both the semi-synthetic and real-world settings that complex narratives (with colliders and forks) perform worse than simple narratives that have a causal chain graph as the ground truth. This gap ,while clear and noticeable, isn’t as stark as failure from parametric knowledge conflict (Table 1) or topological ordering (Figure 4). We do furthermore note that this is one area where extracting an explicit causal graph does not seem to significantly improve performance.
5 Discussion
Our work takes initial strides towards examining the success and failure of LLMs to reason causally on narratives that express causal events. We focus on two questions of key importance in causality: (1) Does one event cause another? (2) Can the LLM extract the causal graph from the narrative. We find three significant failure modes of LLM reasoning by conducting experiments in carefully controlled synthetic, semi-synthetic and real-world settings: Firstly, we find that LLMs rely heavily on topological ordering, performing well when the ordering of events in the narratives matches that of the ordering of the underlying causal graph. Secondly, we find that LLMs rely on their parametric knowledge as a shortcut to infer causal relations. Finally, we examine the role of causal structure complexity, finding that LLM accuracy degrades as the narrative length increases. Furthermore, LLMs perform slightly worse on reasoning when narratives contain structures such as colliders and forks. Beyond these failure modes, we show that more reliable causal reasoning can be elicited by prompting the LLM to explicitly identify the causal graph.
5.1 Limitations and Future Works
One limitation of our work is that there are other forms of causal reasoning that we did not test for in the narratives. This motivates many potential directions for future work. For example, it could be interesting to ask the LLM to reason about counterfactual cases. Our analysis also has implications for algorithmic interventions to improve causal reasoning. The failure modes we identify in this paper could inform the design of targeted synthetic tasks to use in finetuning for improved causal reasoning. Additionally, our findings on the benefits of extracting a causal graph can inform prompt engineering efforts to elicit reliable causal reasoning from language models. We believe investigating both directions represents an exciting direction for future work.
  • Cobbe et al. (2021) Karl Cobbe, Vineet Kosaraju, Mohammad Bavarian, Mark Chen, Heewoo Jun, Lukasz Kaiser, Matthias Plappert, Jerry Tworek, Jacob Hilton, Reiichiro Nakano, Christopher Hesse, and John Schulman. Training verifiers to solve math word problems, 2021. URL https://arxiv.org/abs/2110.14168.
  • Wei et al. (2022) Jason Wei, Xuezhi Wang, Dale Schuurmans, Maarten Bosma, Fei Xia, Ed Chi, Quoc V Le, Denny Zhou, et al. Chain-of-thought prompting elicits reasoning in large language models. Neural Information Processing Systems, 2022.
  • Heindorf et al. (2020) Stefan Heindorf, Yan Scholten, Henning Wachsmuth, Axel-Cyrille Ngonga Ngomo, and Martin Potthast. Causenet: Towards a causality graph extracted from the web. In ACM international conference on Information & Knowledge Management, 2020.
  • Jin et al. (2023) Zhijing Jin, Yuen Chen, Felix Leeb, Luigi Gresele, Ojasv Kamal, Zhiheng LYU, Kevin Blin, Fernando Gonzalez Adauto, Max Kleiman-Weiner, Mrinmaya Sachan, and Bernhard Schölkopf. CLadder: A benchmark to assess causal reasoning capabilities of language models. In Thirty-seventh Conference on Neural Information Processing Systems, 2023. URL https://openreview.net/forum?id=e2wtjx0Yqu.
  • Tan et al. (2022) Fiona Anting Tan, Ali Hürriyetoğlu, Tommaso Caselli, Nelleke Oostdijk, Tadashi Nomoto, Hansi Hettiarachchi, Iqra Ameer, Onur Uca, Farhana Ferdousi Liza, and Tiancheng Hu. The causal news corpus: Annotating causal relations in event sentences from news. In Language Resources and Evaluation Conference. European Language Resources Association, 2022. URL https://aclanthology.org/2022.lrec-1.246.
  • Joshi et al. (2024a) Nitish Joshi, Abulhair Saparov, Yixin Wang, and He He. Llms are prone to fallacies in causal inference, 2024a. URL https://arxiv.org/abs/2406.12158.
  • Jin et al. (2024) Zhijing Jin, Jiarui Liu, Zhiheng LYU, Spencer Poff, Mrinmaya Sachan, Rada Mihalcea, Mona T. Diab, and Bernhard Schölkopf. Can large language models infer causation from correlation? In The Twelfth International Conference on Learning Representations, 2024. URL https://openreview.net/forum?id=vqIH0ObdqL.
  • Gordon et al. (2012) Andrew Gordon, Zornitsa Kozareva, and Melissa Roemmele. SemEval-2012 task 7: Choice of plausible alternatives: An evaluation of commonsense causal reasoning. In Eneko Agirre, Johan Bos, Mona Diab, Suresh Manandhar, Yuval Marton, and Deniz Yuret, editors, *SEM 2012: The First Joint Conference on Lexical and Computational Semantics – Volume 1: Proceedings of the main conference and the shared task, and Volume 2: Proceedings of the Sixth International Workshop on Semantic Evaluation (SemEval 2012), pages 394–398, Montréal, Canada, 7-8 June 2012. Association for Computational Linguistics. URL https://aclanthology.org/S12-1052/.
  • Joshi et al. (2024b) Abhinav Joshi, Areeb Ahmad, and Ashutosh Modi. Cold: Causal reasoning in closed daily activities, 2024b. URL https://arxiv.org/abs/2411.19500.
  • Ho et al. (2023) Matthew Ho, Aditya Sharma, Justin Chang, Michael Saxon, Sharon Levy, Yujie Lu, and William Yang Wang. Wikiwhy: Answering and explaining cause-and-effect questions. In The Eleventh International Conference on Learning Representations, 2023. URL https://openreview.net/forum?id=vaxnu-Utr4l.
  • Zhang et al. (2023) Cheng Zhang, Stefan Bauer, Paul Bennett, Jiangfeng Gao, Wenbo Gong, Agrin Hilmkil, Joel Jennings, Chao Ma, Tom Minka, Nick Pawlowski, et al. Understanding causality with large language models: Feasibility and opportunities. arXiv preprint arXiv:2304.05524, 2023.
  • Wang et al. (2023) Zhaowei Wang, Quyet V. Do, Hongming Zhang, Jiayao Zhang, Weiqi Wang, Tianqing Fang, Yangqiu Song, Ginny Wong, and Simon See. COLA: Contextualized commonsense causal reasoning from the causal inference perspective. In Anna Rogers, Jordan Boyd-Graber, and Naoaki Okazaki, editors, Proceedings of the 61st Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), pages 5253–5271, Toronto, Canada, July 2023. Association for Computational Linguistics. doi: 10.18653/v1/2023.acl-long.288. URL https://aclanthology.org/2023.acl-long.288/.
  • Ashwani et al. (2024) Swagata Ashwani, Kshiteesh Hegde, Nishith Reddy Mannuru, Mayank Jindal, Dushyant Singh Sengar, Krishna Chaitanya Rao Kathala, Dishant Banga, Vinija Jain, and Aman Chadha. Cause and effect: Can large language models truly understand causality?, 2024. URL https://arxiv.org/abs/2402.18139.
  • Zečević et al. (2023) Matej Zečević, Moritz Willig, Devendra Singh Dhami, and Kristian Kersting. Causal parrots: Large language models may talk causality but are not causal, 2023. URL https://arxiv.org/abs/2308.13067.
  • Frohberg and Binder (2022) Jörg Frohberg and Frank Binder. Crass: A novel data set and benchmark to test counterfactual reasoning of large language models, 2022. URL https://arxiv.org/abs/2112.11941.
  • Li et al. (2023) Jiaxuan Li, Lang Yu, and Allyson Ettinger. Counterfactual reasoning: Testing language models’ understanding of hypothetical scenarios, 2023. URL https://arxiv.org/abs/2305.16572.
  • Gao et al. (2023) Jinglong Gao, Xiao Ding, Bing Qin, and Ting Liu. Is chatgpt a good causal reasoner? a comprehensive evaluation, 2023. URL https://arxiv.org/abs/2305.07375.
  • Kıcıman et al. (2024) Emre Kıcıman, Robert Ness, Amit Sharma, and Chenhao Tan. Causal reasoning and large language models: Opening a new frontier for causality, 2024. URL https://arxiv.org/abs/2305.00050.
  • Ammanabrolu et al. (2020) Prithviraj Ammanabrolu, Wesley Cheung, William Broniec, and Mark O. Riedl. Automated storytelling via causal, commonsense plot ordering, 2020. URL https://arxiv.org/abs/2009.00829.
  • Tian et al. (2021) Yufei Tian, Arvind krishna Sridhar, and Nanyun Peng. Hypogen: Hyperbole generation with commonsense and counterfactual knowledge, 2021. URL https://arxiv.org/abs/2109.05097.
  • Li et al. (2022) Bryan Li, Lara J. Martin, and Chris Callison-Burch. $\rm{C{\small IS}}^{2}$: A simplified commonsense inference evaluation for story prose, 2022. URL https://arxiv.org/abs/2202.07880.
  • OpenAI et al. (2024) OpenAI, Josh Achiam, Steven Adler, Sandhini Agarwal, Lama Ahmad, Ilge Akkaya, Florencia Leoni Aleman, Diogo Almeida, Janko Altenschmidt, Sam Altman, Shyamal Anadkat, Red Avila, Igor Babuschkin, Suchir Balaji, Valerie Balcom, Paul Baltescu, Haiming Bao, Mohammad Bavarian, Jeff Belgum, Irwan Bello, Jake Berdine, Gabriel Bernadett-Shapiro, Christopher Berner, Lenny Bogdonoff, Oleg Boiko, Madelaine Boyd, Anna-Luisa Brakman, Greg Brockman, Tim Brooks, Miles Brundage, Kevin Button, Trevor Cai, Rosie Campbell, Andrew Cann, Brittany Carey, Chelsea Carlson, Rory Carmichael, Brooke Chan, Che Chang, Fotis Chantzis, Derek Chen, Sully Chen, Ruby Chen, Jason Chen, Mark Chen, Ben Chess, Chester Cho, Casey Chu, Hyung Won Chung, Dave Cummings, Jeremiah Currier, Yunxing Dai, Cory Decareaux, Thomas Degry, Noah Deutsch, Damien Deville, Arka Dhar, David Dohan, Steve Dowling, Sheila Dunning, Adrien Ecoffet, Atty Eleti, Tyna Eloundou, David Farhi, Liam Fedus, Niko Felix, Simón Posada Fishman, Juston Forte, Isabella Fulford, Leo Gao, Elie Georges, Christian Gibson, Vik Goel, Tarun Gogineni, Gabriel Goh, Rapha Gontijo-Lopes, Jonathan Gordon, Morgan Grafstein, Scott Gray, Ryan Greene, Joshua Gross, Shixiang Shane Gu, Yufei Guo, Chris Hallacy, Jesse Han, Jeff Harris, Yuchen He, Mike Heaton, Johannes Heidecke, Chris Hesse, Alan Hickey, Wade Hickey, Peter Hoeschele, Brandon Houghton, Kenny Hsu, Shengli Hu, Xin Hu, Joost Huizinga, Shantanu Jain, Shawn Jain, Joanne Jang, Angela Jiang, Roger Jiang, Haozhun Jin, Denny Jin, Shino Jomoto, Billie Jonn, Heewoo Jun, Tomer Kaftan, Łukasz Kaiser, Ali Kamali, Ingmar Kanitscheider, Nitish Shirish Keskar, Tabarak Khan, Logan Kilpatrick, Jong Wook Kim, Christina Kim, Yongjik Kim, Jan Hendrik Kirchner, Jamie Kiros, Matt Knight, Daniel Kokotajlo, Łukasz Kondraciuk, Andrew Kondrich, Aris Konstantinidis, Kyle Kosic, Gretchen Krueger, Vishal Kuo, Michael Lampe, Ikai Lan, Teddy Lee, Jan Leike, Jade Leung, Daniel Levy, Chak Ming Li, Rachel Lim, Molly Lin, Stephanie Lin, Mateusz Litwin, Theresa Lopez, Ryan Lowe, Patricia Lue, Anna Makanju, Kim Malfacini, Sam Manning, Todor Markov, Yaniv Markovski, Bianca Martin, Katie Mayer, Andrew Mayne, Bob McGrew, Scott Mayer McKinney, Christine McLeavey, Paul McMillan, Jake McNeil, David Medina, Aalok Mehta, Jacob Menick, Luke Metz, Andrey Mishchenko, Pamela Mishkin, Vinnie Monaco, Evan Morikawa, Daniel Mossing, Tong Mu, Mira Murati, Oleg Murk, David Mély, Ashvin Nair, Reiichiro Nakano, Rajeev Nayak, Arvind Neelakantan, Richard Ngo, Hyeonwoo Noh, Long Ouyang, Cullen O’Keefe, Jakub Pachocki, Alex Paino, Joe Palermo, Ashley Pantuliano, Giambattista Parascandolo, Joel Parish, Emy Parparita, Alex Passos, Mikhail Pavlov, Andrew Peng, Adam Perelman, Filipe de Avila Belbute Peres, Michael Petrov, Henrique Ponde de Oliveira Pinto, Michael, Pokorny, Michelle Pokrass, Vitchyr H. Pong, Tolly Powell, Alethea Power, Boris Power, Elizabeth Proehl, Raul Puri, Alec Radford, Jack Rae, Aditya Ramesh, Cameron Raymond, Francis Real, Kendra Rimbach, Carl Ross, Bob Rotsted, Henri Roussez, Nick Ryder, Mario Saltarelli, Ted Sanders, Shibani Santurkar, Girish Sastry, Heather Schmidt, David Schnurr, John Schulman, Daniel Selsam, Kyla Sheppard, Toki Sherbakov, Jessica Shieh, Sarah Shoker, Pranav Shyam, Szymon Sidor, Eric Sigler, Maddie Simens, Jordan Sitkin, Katarina Slama, Ian Sohl, Benjamin Sokolowsky, Yang Song, Natalie Staudacher, Felipe Petroski Such, Natalie Summers, Ilya Sutskever, Jie Tang, Nikolas Tezak, Madeleine B. Thompson, Phil Tillet, Amin Tootoonchian, Elizabeth Tseng, Preston Tuggle, Nick Turley, Jerry Tworek, Juan Felipe Cerón Uribe, Andrea Vallone, Arun Vijayvergiya, Chelsea Voss, Carroll Wainwright, Justin Jay Wang, Alvin Wang, Ben Wang, Jonathan Ward, Jason Wei, CJ Weinmann, Akila Welihinda, Peter Welinder, Jiayi Weng, Lilian Weng, Matt Wiethoff, Dave Willner, Clemens Winter, Samuel Wolrich, Hannah Wong, Lauren Workman, Sherwin Wu, Jeff Wu, Michael Wu, Kai Xiao, Tao Xu, Sarah Yoo, Kevin Yu, Qiming Yuan, Wojciech Zaremba, Rowan Zellers, Chong Zhang, Marvin Zhang, Shengjia Zhao, Tianhao Zheng, Juntang Zhuang, William Zhuk, and Barret Zoph. Gpt-4 technical report, 2024. URL https://arxiv.org/abs/2303.08774.
  • Anthropic (2024) Anthropic. Claude 3.5 sonnet. 2024. Available at: https://www.anthropic.com/news/claude-3-5-sonnet.
  • Grattafiori et al. (2024) Aaron Grattafiori, Abhimanyu Dubey, Abhinav Jauhri, Abhinav Pandey, Abhishek Kadian, Ahmad Al-Dahle, Aiesha Letman, Akhil Mathur, Alan Schelten, Alex Vaughan, Amy Yang, Angela Fan, Anirudh Goyal, Anthony Hartshorn, Aobo Yang, Archi Mitra, Archie Sravankumar, Artem Korenev, Arthur Hinsvark, Arun Rao, Aston Zhang, Aurelien Rodriguez, Austen Gregerson, Ava Spataru, Baptiste Roziere, Bethany Biron, Binh Tang, Bobbie Chern, Charlotte Caucheteux, Chaya Nayak, Chloe Bi, Chris Marra, Chris McConnell, Christian Keller, Christophe Touret, Chunyang Wu, Corinne Wong, Cristian Canton Ferrer, Cyrus Nikolaidis, Damien Allonsius, Daniel Song, Danielle Pintz, Danny Livshits, Danny Wyatt, David Esiobu, Dhruv Choudhary, Dhruv Mahajan, Diego Garcia-Olano, Diego Perino, Dieuwke Hupkes, Egor Lakomkin, Ehab AlBadawy, Elina Lobanova, Emily Dinan, Eric Michael Smith, Filip Radenovic, Francisco Guzmán, Frank Zhang, Gabriel Synnaeve, Gabrielle Lee, Georgia Lewis Anderson, Govind Thattai, Graeme Nail, Gregoire Mialon, Guan Pang, Guillem Cucurell, Hailey Nguyen, Hannah Korevaar, Hu Xu, Hugo Touvron, Iliyan Zarov, Imanol Arrieta Ibarra, Isabel Kloumann, Ishan Misra, Ivan Evtimov, Jack Zhang, Jade Copet, Jaewon Lee, Jan Geffert, Jana Vranes, Jason Park, Jay Mahadeokar, Jeet Shah, Jelmer van der Linde, Jennifer Billock, Jenny Hong, Jenya Lee, Jeremy Fu, Jianfeng Chi, Jianyu Huang, Jiawen Liu, Jie Wang, Jiecao Yu, Joanna Bitton, Joe Spisak, Jongsoo Park, Joseph Rocca, Joshua Johnstun, Joshua Saxe, Junteng Jia, Kalyan Vasuden Alwala, Karthik Prasad, Kartikeya Upasani, Kate Plawiak, Ke Li, Kenneth Heafield, Kevin Stone, Khalid El-Arini, Krithika Iyer, Kshitiz Malik, Kuenley Chiu, Kunal Bhalla, Kushal Lakhotia, Lauren Rantala-Yeary, Laurens van der Maaten, Lawrence Chen, Liang Tan, Liz Jenkins, Louis Martin, Lovish Madaan, Lubo Malo, Lukas Blecher, Lukas Landzaat, Luke de Oliveira, Madeline Muzzi, Mahesh Pasupuleti, Mannat Singh, Manohar Paluri, Marcin Kardas, Maria Tsimpoukelli, Mathew Oldham, Mathieu Rita, Maya Pavlova, Melanie Kambadur, Mike Lewis, Min Si, Mitesh Kumar Singh, Mona Hassan, Naman Goyal, Narjes Torabi, Nikolay Bashlykov, Nikolay Bogoychev, Niladri Chatterji, Ning Zhang, Olivier Duchenne, Onur Çelebi, Patrick Alrassy, Pengchuan Zhang, Pengwei Li, Petar Vasic, Peter Weng, Prajjwal Bhargava, Pratik Dubal, Praveen Krishnan, Punit Singh Koura, Puxin Xu, Qing He, Qingxiao Dong, Ragavan Srinivasan, Raj Ganapathy, Ramon Calderer, Ricardo Silveira Cabral, Robert Stojnic, Roberta Raileanu, Rohan Maheswari, Rohit Girdhar, Rohit Patel, Romain Sauvestre, Ronnie Polidoro, Roshan Sumbaly, Ross Taylor, Ruan Silva, Rui Hou, Rui Wang, Saghar Hosseini, Sahana Chennabasappa, Sanjay Singh, Sean Bell, Seohyun Sonia Kim, Sergey Edunov, Shaoliang Nie, Sharan Narang, Sharath Raparthy, Sheng Shen, Shengye Wan, Shruti Bhosale, Shun Zhang, Simon Vandenhende, Soumya Batra, Spencer Whitman, Sten Sootla, Stephane Collot, Suchin Gururangan, Sydney Borodinsky, Tamar Herman, Tara Fowler, Tarek Sheasha, Thomas Georgiou, Thomas Scialom, Tobias Speckbacher, Todor Mihaylov, Tong Xiao, Ujjwal Karn, Vedanuj Goswami, Vibhor Gupta, Vignesh Ramanathan, Viktor Kerkez, Vincent Gonguet, Virginie Do, Vish Vogeti, Vítor Albiero, Vladan Petrovic, Weiwei Chu, Wenhan Xiong, Wenyin Fu, Whitney Meers, Xavier Martinet, Xiaodong Wang, Xiaofang Wang, Xiaoqing Ellen Tan, Xide Xia, Xinfeng Xie, Xuchao Jia, Xuewei Wang, Yaelle Goldschlag, Yashesh Gaur, Yasmine Babaei, Yi Wen, Yiwen Song, Yuchen Zhang, Yue Li, Yuning Mao, Zacharie Delpierre Coudert, Zheng Yan, Zhengxing Chen, Zoe Papakipos, Aaditya Singh, Aayushi Srivastava, Abha Jain, Adam Kelsey, Adam Shajnfeld, Adithya Gangidi, Adolfo Victoria, Ahuva Goldstand, Ajay Menon, Ajay Sharma, Alex Boesenberg, Alexei Baevski, Allie Feinstein, Amanda Kallet, Amit Sangani, Amos Teo, Anam Yunus, Andrei Lupu, Andres Alvarado, Andrew Caples, Andrew Gu, Andrew Ho, Andrew Poulton, Andrew Ryan, Ankit Ramchandani, Annie Dong, Annie Franco, Anuj Goyal, Aparajita Saraf, Arkabandhu Chowdhury, Ashley Gabriel, Ashwin Bharambe, Assaf Eisenman, Azadeh Yazdan, Beau James, Ben Maurer, Benjamin Leonhardi, Bernie Huang, Beth Loyd, Beto De Paola, Bhargavi Paranjape, Bing Liu, Bo Wu, Boyu Ni, Braden Hancock, Bram Wasti, Brandon Spence, Brani Stojkovic, Brian Gamido, Britt Montalvo, Carl Parker, Carly Burton, Catalina Mejia, Ce Liu, Changhan Wang, Changkyu Kim, Chao Zhou, Chester Hu, Ching-Hsiang Chu, Chris Cai, Chris Tindal, Christoph Feichtenhofer, Cynthia Gao, Damon Civin, Dana Beaty, Daniel Kreymer, Daniel Li, David Adkins, David Xu, Davide Testuggine, Delia David, Devi Parikh, Diana Liskovich, Didem Foss, Dingkang Wang, Duc Le, Dustin Holland, Edward Dowling, Eissa Jamil, Elaine Montgomery, Eleonora Presani, Emily Hahn, Emily Wood, Eric-Tuan Le, Erik Brinkman, Esteban Arcaute, Evan Dunbar, Evan Smothers, Fei Sun, Felix Kreuk, Feng Tian, Filippos Kokkinos, Firat Ozgenel, Francesco Caggioni, Frank Kanayet, Frank Seide, Gabriela Medina Florez, Gabriella Schwarz, Gada Badeer, Georgia Swee, Gil Halpern, Grant Herman, Grigory Sizov, Guangyi, Zhang, Guna Lakshminarayanan, Hakan Inan, Hamid Shojanazeri, Han Zou, Hannah Wang, Hanwen Zha, Haroun Habeeb, Harrison Rudolph, Helen Suk, Henry Aspegren, Hunter Goldman, Hongyuan Zhan, Ibrahim Damlaj, Igor Molybog, Igor Tufanov, Ilias Leontiadis, Irina-Elena Veliche, Itai Gat, Jake Weissman, James Geboski, James Kohli, Janice Lam, Japhet Asher, Jean-Baptiste Gaya, Jeff Marcus, Jeff Tang, Jennifer Chan, Jenny Zhen, Jeremy Reizenstein, Jeremy Teboul, Jessica Zhong, Jian Jin, Jingyi Yang, Joe Cummings, Jon Carvill, Jon Shepard, Jonathan McPhie, Jonathan Torres, Josh Ginsburg, Junjie Wang, Kai Wu, Kam Hou U, Karan Saxena, Kartikay Khandelwal, Katayoun Zand, Kathy Matosich, Kaushik Veeraraghavan, Kelly Michelena, Keqian Li, Kiran Jagadeesh, Kun Huang, Kunal Chawla, Kyle Huang, Lailin Chen, Lakshya Garg, Lavender A, Leandro Silva, Lee Bell, Lei Zhang, Liangpeng Guo, Licheng Yu, Liron Moshkovich, Luca Wehrstedt, Madian Khabsa, Manav Avalani, Manish Bhatt, Martynas Mankus, Matan Hasson, Matthew Lennie, Matthias Reso, Maxim Groshev, Maxim Naumov, Maya Lathi, Meghan Keneally, Miao Liu, Michael L. Seltzer, Michal Valko, Michelle Restrepo, Mihir Patel, Mik Vyatskov, Mikayel Samvelyan, Mike Clark, Mike Macey, Mike Wang, Miquel Jubert Hermoso, Mo Metanat, Mohammad Rastegari, Munish Bansal, Nandhini Santhanam, Natascha Parks, Natasha White, Navyata Bawa, Nayan Singhal, Nick Egebo, Nicolas Usunier, Nikhil Mehta, Nikolay Pavlovich Laptev, Ning Dong, Norman Cheng, Oleg Chernoguz, Olivia Hart, Omkar Salpekar, Ozlem Kalinli, Parkin Kent, Parth Parekh, Paul Saab, Pavan Balaji, Pedro Rittner, Philip Bontrager, Pierre Roux, Piotr Dollar, Polina Zvyagina, Prashant Ratanchandani, Pritish Yuvraj, Qian Liang, Rachad Alao, Rachel Rodriguez, Rafi Ayub, Raghotham Murthy, Raghu Nayani, Rahul Mitra, Rangaprabhu Parthasarathy, Raymond Li, Rebekkah Hogan, Robin Battey, Rocky Wang, Russ Howes, Ruty Rinott, Sachin Mehta, Sachin Siby, Sai Jayesh Bondu, Samyak Datta, Sara Chugh, Sara Hunt, Sargun Dhillon, Sasha Sidorov, Satadru Pan, Saurabh Mahajan, Saurabh Verma, Seiji Yamamoto, Sharadh Ramaswamy, Shaun Lindsay, Shaun Lindsay, Sheng Feng, Shenghao Lin, Shengxin Cindy Zha, Shishir Patil, Shiva Shankar, Shuqiang Zhang, Shuqiang Zhang, Sinong Wang, Sneha Agarwal, Soji Sajuyigbe, Soumith Chintala, Stephanie Max, Stephen Chen, Steve Kehoe, Steve Satterfield, Sudarshan Govindaprasad, Sumit Gupta, Summer Deng, Sungmin Cho, Sunny Virk, Suraj Subramanian, Sy Choudhury, Sydney Goldman, Tal Remez, Tamar Glaser, Tamara Best, Thilo Koehler, Thomas Robinson, Tianhe Li, Tianjun Zhang, Tim Matthews, Timothy Chou, Tzook Shaked, Varun Vontimitta, Victoria Ajayi, Victoria Montanez, Vijai Mohan, Vinay Satish Kumar, Vishal Mangla, Vlad Ionescu, Vlad Poenaru, Vlad Tiberiu Mihailescu, Vladimir Ivanov, Wei Li, Wenchen Wang, Wenwen Jiang, Wes Bouaziz, Will Constable, Xiaocheng Tang, Xiaojian Wu, Xiaolan Wang, Xilun Wu, Xinbo Gao, Yaniv Kleinman, Yanjun Chen, Ye Hu, Ye Jia, Ye Qi, Yenda Li, Yilin Zhang, Ying Zhang, Yossi Adi, Youngjin Nam, Yu, Wang, Yu Zhao, Yuchen Hao, Yundi Qian, Yunlu Li, Yuzi He, Zach Rait, Zachary DeVito, Zef Rosnbrick, Zhaoduo Wen, Zhenyu Yang, Zhiwei Zhao, and Zhiyu Ma. The llama 3 herd of models, 2024. URL https://arxiv.org/abs/2407.21783.
  • Dettki et al. (2025) Hanna M. Dettki, Brenden M. Lake, Charley M. Wu, and Bob Rehder. Do large language models reason causally like us? even better?, 2025. URL https://arxiv.org/abs/2502.10215.
  • Callan (2012) Jamie Callan. The lemur project and its clueweb12 dataset. In Invited talk at the SIGIR 2012 Workshop on Open-Source Information Retrieval, 2012.
Technical Appendices and Supplementary Material
Appendix A Synthetic Data Experiments
A.1 Selected Synthetic Prompts
We use an LLM to generate the events E. From the events, we create a ground truth causal graph G which is used to structure and inform the narrative sequence and causality. N is the corresponding narrative created by the LLM from G. To evaluate the LLM’s performance, we extract a causal graph, G’, from the narrative N as produced by the LLM, and compare it with the ground truth causal graph G. In this context, n refers to the number of events to generate, while A and B represent pairs of events queried for causal relationships. The task then becomes assessing whether event A causes event B . All prompts, data processing steps, and results are included in the attached code.
A.1.1 Topological Experiment - Generating Random Events (E)
“generate n random distinct events"
A.1.2 Parametric Experiment -Generating a Pair of Causal Events (E)
“generate a pair of events that cause each other. generate an event that causes another event, for example Cancer $\rightarrow$ Death or Obesity $\rightarrow$ Bad Heart Health. Make sure the event generated is not already in $E$
This is repeated as many times as is necessary
A.1.3 Parametric Experiment - Generating a Pair of Anti-Causal Events (E)
“generate a pair of events that are anticausal (an event causing the opposite of the normal effect), for example the first event could be cancer and the second event could be a longer life because in reality, cancer causes a shorter life. Make sure the events generated are not already in $E$."
This is repeated as many times as is necessary
A.1.4 Forward Topological Narrative (N)
“Output a short narrative (use one sentence) that expresses the causal link [E1 $\rightarrow$ E2]. By causal link, we mean that the sentence should convey that E1 directly caused E2. In other words, it should be clear from the narrative that E2 would not have happened had E1 not happened. Ensure that the words [E1, E2] are present in the new sentence and E1 appears before E2. Only output the new sentence."
Repeat for all causal/anti-causal links
A.1.5 Reverse Topological Narrative (N)
“Output a short narrative (use one sentence) that expresses the causal link [E1 $\rightarrow$ E2]. By causal link, we mean that the sentence should convey that E1 directly caused E2. In other words, it should be clear from the narrative that E2 would not have happened had E1 not happened. Ensure that the words [E1, E2] are present in the new sentence and E2 appears before E1. Only output the new sentence."
Repeat for all causal/anti-causal links
A.1.6 Standard Prompt
“Use this narrative N as context. Did A cause B? Output your answer with $<answer>Yes/No</answer>$. The cause can be direct or indirect."
A.1.7 Chain of Thought Prompt
“Use this narrative N as context. Did A cause B? Do step by step reasoning. Then output your answer with $<answer>Yes/No</answer>$. The cause can be direct or indirect."
A.1.8 In-Context Prompt
“Use this narrative N as context. Did A cause B? Output your answer with $<answer>Yes/No</answer>$. The cause can be direct or indirect. An example narrative would be: Rains leads to plants growing. This then causes increased oxygen in the atmosphere. A potential question would be does rain cause increased oxygen in the atmosphere? The answer would be Yes. Another example narrative would be: Increased oxygen in the atmosphere is because of plants growing. Plants grow because rain provides them essential nutrients. A potential question would be does rain cause increased oxygen in the atmosphere? The answer would be Yes. Another example narrative would be: Rain leads plants to grow. Plants growing causes less oxygen in the atmosphere. A potential question would be does rain cause less oxygen in the atmosphere? The answer would be Yes."
A.1.9 Narrative + Graph Prompt
“Use this narrative N and this causal ordering G’ ((such that each item is a cause of every item after it, for example the first list item is a cause of the third, fourth, fifth items etc)) as context. Did A cause B? Output your answer with $<answer>Yes/No</answer>$. The cause can be direct or indirect."
A.2 Parametric Graph Experiment
Let’s call the graph of parametric knowledge $P$. We then take the odd indexed events (1st, 3rd etc) from $P$ and place them in the first half of the causal ground truth graph $G$ and the even indexed events (2nd, 4th etc) from $P$ in the second half of $G$. This process is shown in Figure 6.
Figure 6: Example illustration (right) is of how $G$, the ground truth causality, is set up.
A.3 Complex Graph Creation
To generate a ground-truth causal graph G with rich structure (colliders, forks, and a spanning chain), for each choice of size $n$ we perform the following algorithm:
  1. 1.
    Node sampling. Draw $n$ distinct events
    $\{E_{1},E_{2},\dots,E_{n}\}\;\subset\;\mathcal{E}$
    uniformly at random without replacement.
  2. 2.
    Determine motif counts. (for $n\geq 4$)
    $k_{\max}\;=\;\bigl{\lfloor}n/2\bigr{\rfloor},\quad k_{\mathrm{tot}}\sim\mathrm{Uniform}\bigl{(}2,\,k_{\max}\bigr{)},$
    $k_{\mathrm{col}}\sim\mathrm{Uniform}\bigl{(}1,\,k_{\mathrm{tot}}-1\bigr{)},\qquad k_{\mathrm{fork}}=k_{\mathrm{tot}}-k_{\mathrm{col}}.$
  3. 3.
    Collider creation. Repeat $k_{\mathrm{col}}$ times:
    1. (a)
      Select two distinct “parent” nodes $p_{1},p_{2}$ from those not yet used in any motif.
    2. (b)
      Select a “child” node $c$ that is neither $p_{1}$ nor $p_{2}$ and not yet used as a child.
    3. (c)
      Add edges
      $p_{1}\;\to\;c\quad\text{and}\quad p_{2}\;\to\;c\,,$
      thereby forming a collider at $c$.
  4. 4.
    Fork creation. Repeat $k_{\mathrm{fork}}$ times:
    1. (a)
      Select a “parent” node $p$ from those not yet used.
    2. (b)
      Select two distinct “child” nodes $c_{1},c_{2}$ from the remaining unused nodes.
    3. (c)
      Add edges
      $p\;\to\;c_{1}\quad\text{and}\quad p\;\to\;c_{2}\,,$
      forming a fork with shared parent $p$.
  5. 5.
    Chain-connect remaining nodes. Let $\mathcal{R}$ be the set of nodes not yet involved in any collider or fork.
    1. (a)
      Order =$\mathcal{R}=\{r_{1},\dots,r_{m}\}$ arbitrarily, then add chain edges
      $r_{1}\to r_{2},\;\;r_{2}\to r_{3},\;\;\dots,\;\;r_{m-1}\to r_{m}.$
    2. (b)
      To ensure the entire graph is connected, choose one node $u$ from among the previously used nodes (if any) and add
      $u\;\to\;r_{1}\,.$
Appendix B Real-world Causal Graphs
B.1 Prompt templates for narrative generation
Recall that we have a ground truth causal chain graph of the form $V_{1}\rightarrow V_{2}\rightarrow\ldots\rightarrow V_{N}$ from CauseNet that we need to verbalize into a coherent narrative. For the semi-synthetic narratives, we use the LLM (GPT-4o) to do so one edge at a time, while ensuring that the newly verbalized edge logically follows the previous one. The following is the prompt template for generating the narratives in the topological order of the graph:
Output a short narrative (use one or two sentences) that expresses the causal link [$V_{i}\rightarrow V_{i+1}$] and logically follows this narrative:
{ Narrative for the previous edge $V_{i-1}\rightarrow V_{i}$}.
Ensure that the combined sentences convey the causal chain [$V_{i-1}\rightarrow V_{i}\rightarrow V_{i+1}$] and that the words [$V_{i}$, $V_{i+1}$] are present. Only output the newly generated narrative.
Similarly, we generate narratives in the reverse topological order of the graph by verbalizing edges in the reverse direction with the following prompt template:
Output a short narrative (use one or two sentences) that expresses the causal link [$V_{i}\rightarrow V_{i+1}$] and logically follows this narrative:
{ Narrative for the previous edge $V_{i+1}\rightarrow V_{i+2}$}.
Ensure that the combined sentences convey the causal chain [$V_{i}\rightarrow V_{i+1}\rightarrow V_{i+2}$] and that the words [$V_{i}$, $V_{i+1}$] are present. Only output the newly generated narrative.
For generating real-world narratives, for each edge $V_{i}\rightarrow V_{j}$, we use the set of sentences from CauseNet. Each edge in CauseNet is linked to multiple sentences from various sources. Picking a sentence for each edge at random and concatenating them does not always lead to sensible narratives. To improve the quality of narratives, we use the following prompt to concatenate sentences for adjacent edges:
Consider the following sentences.
{ Sentence for edge $V_{i}\rightarrow V_{i+1}$. { Sentence for edge $V_{i+1}\rightarrow V_{i+2}$.
Do the sentences logically follow each other and express the causal chain [$V_{i}\rightarrow V_{i+1}\rightarrow V_{i+2}$]? Answer with Yes or No.
For verbalizing narratives in the topological order, for a given graph $V_{1}\rightarrow V_{2}\rightarrow\ldots\rightarrow V_{N}$, we only use sentences such that the above prompt returns Yes for every pair of adjacent edges $V_{i}\rightarrow V_{i+1}\rightarrow V_{i+2}$. This ensures that the narrative as a whole remains coherent and conveys the entire causal chain graph. We use a similar prompting strategy to verbalize narratives in the reverse topological order.
B.2 Eliciting Parametric Knowledge
We ask the LLM “Does $V_{i}$ typically have a causal (indirect or direct) effect on $V_{j}$?” and “Would it be atypical if $V_{i}$ had a (indirect or direct) causal effect on $V_{j}$?”. If the LLM answers “No” and “Yes” to those respective questions, we would consider a causal relationship between $V_{i}$ and $V_{j}$ to contradict the LLM’s prior knowledge that it learned from its pretraining corpora.
B.3 Semi-Synthetic and Real-World Complex Graph Algorithm
Let =$\mathcal{M}=\{(u,v)\}$ be the set of real-world causal edges from CauseNet. For each target size $n\in\{3,\dots,9\}$, we:
  1. 1.
    Load CauseNet.
    =$\mathcal{M}\;=\;\bigl{\{}(u,v)\mid u\!\to\!v\text{ in CauseNet}\bigr{\}}.$
  2. 2.
    Extract collider and fork motifs.
    $\mathsf{Colliders}$ =$=\{(p_{1},p_{2},c)\mid(p_{1},c)\in\mathcal{M},\,(p_{2},c)\in\mathcal{M},\,p_{1}\neq p_{2}\},$
    $\mathsf{Forks}$ =$=\{(r,c_{1},c_{2})\mid(r,c_{1})\in\mathcal{M},\,(r,c_{2})\in\mathcal{M},\,c_{1}\neq c_{2}\}.$
  3. 3.
    Determine motif counts.
    $\text{If }n=3,\quad(k_{\mathrm{col}},k_{\mathrm{fork}})=\begin{cases}(1,0)\quad&\text{w.p.\ }0.5,\\ (0,1)\quad&\text{w.p.\ }0.5.\end{cases}$
    (for $n\geq 4$)
    $k_{\max}\;=\;\bigl{\lfloor}n/2\bigr{\rfloor},\quad k_{\mathrm{tot}}\sim\mathrm{Uniform}\bigl{(}2,\,k_{\max}\bigr{)},$
    $k_{\mathrm{col}}\sim\mathrm{Uniform}\bigl{(}1,\,k_{\mathrm{tot}}-1\bigr{)},\qquad k_{\mathrm{fork}}=k_{\mathrm{tot}}-k_{\mathrm{col}}.$
  4. 4.
    Select motifs.
    • Sample $k_{\mathrm{col}}$ distinct triples from $\mathsf{Colliders}$.
    • Sample $k_{\mathrm{fork}}$ distinct triples from $\mathsf{Forks}$.
    Let $S$ be the union of all nodes appearing in these sampled triples.
  5. 5.
    Pad or trim to size $n$.
    • If $|S|>n$, uniformly subsample $n$ nodes from $S$.
    • If $|S|<n$, add random “seed” nodes (not already in $S$) until $|S|=n$.
  6. 6.
    Build ground-truth edges $\mathcal{G}\subseteq S\times S$.
    1. (a)
      Colliders: for each $(p_{1},p_{2},c)$ chosen, add $p_{1}\to c$ and $p_{2}\to c$.
    2. (b)
      Forks: for each $(r,c_{1},c_{2})$, add $r\to c_{1}$ and $r\to c_{2}$.
    3. (c)
      Chains: for any remaining $(u,v)\in S\times S$ with $(u,v)\in\mathcal{M}$ and neither $u$ nor $v$ used in the above, add $u\to v$ to ensure connectivity.
  7. 7.
    Narrative generation. For each $(u\!\to\!v)\in\mathcal{G}$:
    For the semi-synthetic case - prompt the LLM to generate a sentence linking u to v using the forward topological ordering prompt.
    For the real-world case: Find a causal sentence linking u and v in the Cause-Net database
Appendix C Real-World Complex Graph Creation
C.1 Additional examples of the generated narratives
C.1.1 Semi-synthetic narratives
Below, we present some examples of semi-synthetic narratives in the forward and reverse directions.
The narrative in the forward direction for the chain higher prices $\rightarrow$ reduced demand $\rightarrow$ lower prices:
As higher prices swept through the market, consumers began to tighten their budgets, leading to a noticeable reduction in demand for many goods. As a result of the reduced demand, suppliers were forced to lower prices in order to attract buyers back to the market.
The narrative in the reverse order for the causal chain bankruptcy $\rightarrow$ bad credit $\rightarrow$ rejection $\rightarrow$ anger:
The sting of rejection ignited a fire within her, transforming her hurt into a seething anger that demanded to be felt. Her bad credit had led to the rejection she never saw coming, and now that sting of rejection ignited a fire within her, transforming her hurt into a seething anger that demanded to be felt. Her bankruptcy had left her with bad credit, a shadow that loomed over her every application, and now that sting of rejection ignited a fire within her, transforming her hurt into a seething anger that demanded to be felt.
The narrative in the reverse order for the causal chain pollution $\rightarrow$ climate change $\rightarrow$ extreme weather events $\rightarrow$ natural disasters:
As extreme weather events become more frequent and severe, they increasingly lead to devastating natural disasters that disrupt communities and ecosystems alike. Climate change is driving the rise in extreme weather events, which in turn are causing unprecedented natural disasters that threaten the stability of communities and the health of ecosystems. Pollution is a major contributor to climate change, which is driving the rise in extreme weather events that threaten the stability of communities and the health of ecosystems.
C.1.2 Real-world narratives
Below, we present some examples of real-world narratives in the forward and reverse directions.
The narrative in the forward direction for the chain higher prices $\rightarrow$ reduced demand $\rightarrow$ lower prices:
Higher prices generally lead to reduced demand. Lower prices, caused by reduced demand and increased competition for soybeans and corn, largely contributed to the overall bulk export decline.
The narrative in the reverse order for the causal chain bankruptcy $\rightarrow$ bad credit $\rightarrow$ rejection $\rightarrow$ anger:
Embittered by an abusive upbringing, seething with resentment, irritated by others’ failure to fulfill his or her superior sense of entitlement, and fuelled by anger resulting from rejection, the serial bully displays an obsessive, compulsive and self-gratifying urge to displace their uncontrolled aggression onto others whilst exhibiting an apparent lack of insight into their behavior and its effect on people around them. Bad credit normally leads to rejection but now with bad credit secured loan, you can avail the loan of your choice. For example, if you are applying for a loan, the lender may reject your application on the basis of bad credit caused by bankruptcy.
The narrative in the reverse order for the causal chain pollution $\rightarrow$ climate change $\rightarrow$ extreme weather events $\rightarrow$ natural disasters:
In addition to forced migrations from rising seas, climate change is also increasing extreme weather events causing natural disasters such as cyclonic storms (hurricanes or typhoons), floods and droughts. This is worsened by extreme weather events caused by climate change. This landmark bill would jump start the economy by creating millions of new clean energy jobs, increase national security by reducing dependence on foreign oil, and preserve the planet by reducing the pollution that causes climate change.
C.2 Prompt templates for assessing causal reasoning
We use the following template for the Direct prompting strategy:
Consider the following hypothetical narrative.
{narrative}
According to the hypothetical narrative, does {cause} have a (direct or indirect) causal effect on {effect}? Answer in Yes/No.
We use the following template for the Chain-of-Though (CoT) prompting strategy:
Consider the following hypothetical narrative.
{narrative}
According to the hypothetical narrative, does {cause} have a (direct or indirect) causal effect on {effect}? Think step-by-step and end your answer with <answer>Yes/No</answer>.
We use the following template to extract a chain graph from the narrative:
Consider the following hypothetical narrative.
{narrative}
[…]
C.3 Necessary Compute
No pretraining was done so no GPUs were needed. We used cloud based API calls to pre-trained models like ChatGPT, Anthropic and Llama. We estimate that for the synthetic portion, our API calls to ChatGPT, Anthropic and LLama took 10 hours each. For the semi-synthetic and real-world portion, we had roughly 10 hours of API calls for ChatGPT and Llama each. So in total, roughly 50 hours of API usage. As the majority of the computational burden fell on cloud based API calls, no significant CPU resources are required either.
Appendix D Additional Results - Synthetic Data
D.1 Forward vs Reverse Experiments Anthropic and LLama
(a) Anthropic Claude 3.5 Sonnet
(b) LLama 3.1 8B
[…]
D.2 Causal Vs Anti-Causal Experiments Anthropic and LLama
(a) Anthropic Claude 3.5 Sonnet
(b) LLama 3.1 8B
[…]
D.3 Complex vs Simple Graphs Anthropic and LLama
(a) Anthropic Claude 3.5 Sonnet
(b) LLama 3.1 8B
[…]
Appendix E Additional results - Semi-Synthetic and Real World Data
E.1 Forward vs Reverse LLama
Figure 10: (LLama 3.1 8B) The accuracy of various prompting strategies (error bars denote 95% CIs) in the Semi-Synthetic and Real-World Regimes using CauseNet. We observe that the accuracy is lower in the reverse direction .
E.2 Parametric Experiment LLama
Standard CoT Graph
Semi-synthetic
Without Conflict 88.4 83.7 99.5
With Conflict 61.4 57.9 98.2
Real-world
Without Conflict 81.6 79.2 95.1
With Conflict 48.8 49.9 93.2
Table 2: (LLama 3.1 8B) The average accuracy across different narratives with the three prompting strategies partitioned by whether the cause-effect pairs conflict with the LLM’s parametric knowledge (we omit the 95% CIs as they are smaller than $0.3$).
E.3 Simple vs Complex LLama
[…]
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: Yes all claims made are backed up by rigorous empirical evaluations on multiple LLMs in synthetic, semi-synthetic and real-world settings.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: Yes limitations are discussed in the Discussion section including the fact that we only assess one type of causal reasoning (and not counterfactuals etc).
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: We have no theoretical results.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: Yes all prompts used are in the Appendix along with example narratives, and the full set of data and narratives used can be found in the code.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We provide the highly commented code/prompts/narratives with the submission so that all specified details can be seen. This code will also be made public if the paper is accepted.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: Yes, we provide exact details on all prompts used to generate responses in the appendix as well as provide psuedocode readers can follow to intuitively understand the algorithms used such as the one to generate complex graphs. Furthermore, all specific details such as generated narratives and model parameters are provided in the code that is attached and that will be made public.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: Confidence intervals are provided in the form of bootstrapped CIs in plots and it is specified they are confidence intervals for the mean.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: Yes total time used for experiments is discussed in the Appendix, as well as the fact that no GPU usage or significant CPU usage was required (as we used cloud based APIs).
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We have no violations of the code.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We discuss societal benefits if LLMs are able to better understand causal content in narratives. There is no foreseeable negative consequences on society of our work.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: Our paper does not create models/datasets that have high risk of misuse.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We only use existing data that is freely available to researchers, with the source cited.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: Our code contains comments and instructions for running any necessary models.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: No human subjects data was collected
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: No human subjects data was collected
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: No LLMs were used to produce any part of this paper.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.